OpenStudy (anonymous):

Ellen bought gifts for her friends, and each gift cost $7. Let y represent the total cost and x represent the number of gifts. Write an equation to represent the proportional relationship between the total cost and the number of gifts she bought.

OpenStudy (anonymous):

do you know why?

OpenStudy (anonymous):

Tell me if it is correct or not

OpenStudy (anonymous):

I know if it is, but you tell me

OpenStudy (blackstreet23):

ohh yeah that makes sense

OpenStudy (anonymous):

write the equation you think it is

OpenStudy (anonymous):

and compare it to mine

OpenStudy (anonymous):

and I'll tell you if it is right or not

OpenStudy (anonymous):

x*7=y, or it can also be 7x=y

OpenStudy (anonymous):

i dont get it

OpenStudy (anonymous):

x is the number of gifts

OpenStudy (anonymous):

7 is really the $7 in the question

OpenStudy (anonymous):

so, the number of gifts times 7 is the cost of all of the gifts

OpenStudy (anonymous):

do you get it now?

OpenStudy (anonymous):

the cost of all of the gifts is y

OpenStudy (anonymous):

im going to send oppsions

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

send them

OpenStudy (anonymous):

y=7x y=1/7x y=6x y=1/6x

OpenStudy (anonymous):

y=7x

OpenStudy (anonymous):

x*7=y is the same as y=7x

OpenStudy (anonymous):

do you know why?

OpenStudy (anonymous):

thank you very much

OpenStudy (anonymous):

okay :), but do you know why?

OpenStudy (anonymous):

x * 7=y